ADNAN MUHAMMED

Grade 12,

A 8000 kg engine pulls a train of 5 wagons ,each of 2000 kg along a horizontal track.If the engine exerts a force of 40000 N and the track offers a friction force of 5000 N,then calculate: (a)the net accelerating force; (b)the acceleration of the train ;and (c)the force of wagon 1 on wagon 2.

8000 kg engine pulls a train of  5 wagons ,each of 2000 kg along a horizontal track.If the engine exerts a force of 40000 N and the track offers a friction force of 5000 N,then calculate:


(a)the net accelerating force;


(b)the acceleration of the train ;and


(c)the force of wagon 1 on wagon 2. 

Grade:Upto college level

5 Answers

Chetan Mandayam Nayakar
12 years ago

a)35000 N

b)35000/(5*2000+8000)='a'

c) a*8000

Daksh Sammi
6 years ago
) The net accelerating force= Force exerted by engine – Force of fricition= 40000 N – 5000 N = 35000 N(b) The acceleration of the trainWe know that, F = mass x accelerationOr, 35000 N = (mass of engine + mass of 5 wagons) X aOr, 35000 N = (8000 kg + 10000 kg) X aOr, 35000N = 18000 kg X a⇒a=35000N18000 kg=1.944ms−2⇒a=35000N18000 kg=1.944ms-2(c) The force of wagon 1 on wagon 2Since, net accelerating force = 35000 NMass of all 5 wagons = 10000 kgWe know that, F = m x aTherefore, 35000N = 10000 kg x a⇒a=35000N10000 kg=3.5ms−2⇒a=35000N10000 kg=3.5ms-2Therefore, acceleration of wagons =3.5m/s2=3.5m/s2Thus, force of wagon 1 on 2 = mass of four wagons x acceleration⇒F=4×2000 kg×3.5ms−2⇒F=4×2000 kg×3.5ms-2⇒F=8000 kg×3.5ms−2⇒F=8000 kg×3.5ms-2⇒F=28000N⇒F=28000NThus, (a) The net accelerating force = 35000N(b) The acceleration of train = 1.944 ms–2(c) The force of wagon 1 on 2 = 28000N
RItesh Prasad
5 years ago
( a ) The net accelerating force = Force exerted by engine – Force of fricition = 40000  – 5000  = 35000 N
​( b ) Net mass = mass of engine + mass of 5 wagons = 8000 + 5x2000 = 18000kg
Accelaration = net force/net mass = 35000/18000 = 1.944 m/s2
( c ) Force of wagon 1 on 2 = mass of four wagons x acceleration = (4x2000)x1.944 = 15556 N
 
Aradhana Nayak
5 years ago
(a) the net accelerating force is 35000N 
(b)It is given that the force is applied on the wagons by the engine. So the weight of the engine should not be considered
By calculation,the acceleration of the train is equal to3.5m/s^2
 
ankit singh
3 years ago
a ) The net accelerating force = Force exerted by engine – Force of fricition = 40000  – 5000  = 35000 N
​( b ) Net mass = mass of engine + mass of 5 wagons = 8000 + 5x2000 = 18000kg
Accelaration = net force/net mass = 35000/18000 = 1.944 m/s2
( c ) Force of wagon 1 on 2 = mass of four wagons x acceleration = (4x2000)x1.944 = 15556 N